# Generative AI for Physical-Layer Authentication

### v2

@@ Abstract
In recent years, Artificial Intelligence (AI)-driven Physical-Layer Authentication (PLA), which focuses on achieving endogenous security and intelligent identity authentication, has attracted considerable interest. When compared with Discriminative AI (DAI), Generative AI (GAI) offers several advantages, such as fingerprint data augmentation, fingerprint denoising and reconstruction, and protection against adversarial attacks. Inspired by these innovations, this paper provides a systematic exploration of GAI’s integration into PLA frameworks. We commence with a review of representative authentication techniques, emphasizing PLA’s inherent strengths. Following this, we revisit four typical GAI models and contrast the limitations of DAI with the potential of GAI in addressing PLA challenges, including insufficient fingerprint data, environment noises and inferences, perturbations in fingerprint data, and complex tasks. Specifically, we delve into providing GAI-enhanced methods for PLA across the fingerprint collection, model training, and performance optimization phases in detail. Moreover, we present a case study that combines fingerprint extrapolation and Generative Diffusion Model (GDM) to illustrate the superiority of GAI in bolstering the reliability of PLA. Additionally, we outline potential future research directions for GAI-based PLA.
Index Terms:
Generative AI, physical-layer authentication, 6G, identity security.
@@ I Introduction
Traditional Discriminative AI (DAI) utilizes deep neural networks to learn features from large-scale fingerprint datasets. However, these schemes rely on extensively labeled fingerprints for training. Fortunately, the emergence of Generative AI (GAI) has mitigated these limitations. Instead of learning class boundaries, GAI learns the probability distribution of training fingerprints to generate new fingerprint samples [8], which is expected to be applied to Internet of Things (IoT) [1, 9], Industrial IoT (IIoT) [3], Internet of Vehicles (IoV) [10], and satellite communication systems [6]. Nevertheless, research on GAI-based PLA is still in its early phases. The full potential of applying GAI to PLA has yet to be comprehensively investigated. Specifically, the following questions remain unanswered:
  • Q1: Why is GAI suitable for PLA?
  • Q2: What issues can GAI handle in the context of PLA?
  • Q3: How does GAI handle these issues?
To address these challenges, we review representative identity authentication approaches and GAI models, and assess the feasibility of applying GAI to PLA. The primary contributions are outlined as:
  • Firstly, we revisit classic identity authentication techniques and prominent GAI models. Then, we compare the limitations of DAI with the potential of GAI in PLA.
  • We sort out that GAI can solve the issues of PLA from the perspectives of fingerprint collection, model training, and performance optimization.
  • We present a case study to demonstrate the effectiveness of GAI in enhancing authentication performance. Additionally, we envision future research directions for GAI-based PLA.
@@ II-A1 Why Identity Authentication?
  • Preventing Unauthorized Access: The open nature of wireless channels exposes signals to public environments, allowing attackers to impersonate legitimate terminals, steal sensitive information, and even tamper with data streams through Man-in-the-Middle attacks. Identity authentication verifies the authenticity of communicating parties, blocks unauthorized devices at the source, and establishes a trusted communication link [3].
  • Safeguarding Data Security: Enabling encryption without prior authentication may allow attackers to use forged identities to participate in key negotiation processes, resulting in encryption key leakage. Identity authentication acts as a trust anchor for encryption systems, ensuring that keys are exclusively held by legitimate parties, thereby preserving the confidentiality and integrity of data transmission [9].
@@ II-A2 How to Identity Authentication?
Cryptography-based authentication is primarily implemented through cryptographic algorithms in the upper-layers of communication systems; Blockchain-based authentication represents a distributed architecture-based group identification authentication method, with the characteristics of a multi-layer framework and a fault-tolerant mechanism; Physical-Layer Key-based Authentication (PLKA) is realized using physical-layer key, which is generated based on the wireless channel reciprocity; PLA is realized based on physical-layer attributes. The detailed descriptions are shown in Table I.
TABLE I: Descriptions of four representative identity authentication techniques, where cryptography-based authentication algorithms are commercial from 2G to 5G systems, blockchain-based authentication is a distributed framework-based group identity authentication approach, Physical-Layer Key-Based Authentication (PLKA) is achieved leveraging the reciprocity of wireless links between the transmitter and receiver, and Physical-Layer Authentication (PLA) is realized based on channel fingerprints and Radio Frequency (RF) fingerprints.
Techniques Descriptions
Cryptography-Based Authentication • 2G: Mainly applied in two systems: global systems for GSM and CDMA. The former generates signed response through key agreement protocols to realize access authentication, while the latter generates verification value through random number calculation. • 3G: Introducing the EAP-AKA protocol to realize the integrity protection of signaling, voice, and data. • 4G: Adopting the Evolved Packet System AKA (EPS-AKA) protocol, adopting signaling separation architectures of access and non-access layers, and resisting internal attacks through differentiated key distribution strategy, but lacking an effective key leakage response mechanism. • 5G: Realizing security upgrade based on EAP-AKA’ and 5G AKA protocols. Enhancements include introducing sequence numbers to defend against replay attacks and using asymmetric encryption to prevent international user identity information from being intercepted. •  Cons: Reliance on the limited computation power assumptions of attackers, difficulty in defense against replay attacks, and requirement for key generation, distribution and updates.
Blockchain-Based Authentication • distributed architecture-based group identity authentication method, with its characteristics primarily embodied in multi-tier frameworks. • During system operation, each node relies on consensus mechanisms and competitive computing to achieve continuous blockchain updates. When data from a single node is illegally tampered with, the system will automatically suspend the update permission of that node. Thanks to the fault-tolerant properties of distributed ledgers, communication between remaining nodes can still proceed normally. •  Constructing security barriers for user access through asymmetric encryption, while employing PKI to implement identity protections. •  Specifically, data layers ensure information immutability, network layers define terminal authentication mechanisms, consensus layers establish pre-agreed authentication protocols for all terminals, incentive layers motivate active participation in authentication tasks across all terminals, and contract layers contain intelligent identification algorithms and smart contracts for anomaly detection. •  Cons: High delay and low throughput.
PLKA • Be realized using physical-layer key, which is generated based on the inherent physical characteristics of wireless channels, including channel fading and noises. By measuring and analyzing these characteristics, keys closely related to specific links can be generated. • The reciprocity of channels ensures that the transmitter and receiver can generate consistent keys based on the same channel characteristics. • The physical characteristics are highly random and unpredictable, which endows the generated key with strong randomness and difficulty in copying, thus greatly improving the security of the key. • The physical characteristics of wireless channels change with time, environment, and other factors, which provides the possibility of frequently updating the key and further strengthens the security performance. •  The steps of its generation include channel detection, channel quantization, information negotiation, and privacy enhancement.It is highly sensitive to minor changes in input fingerprints. •  Cons: Difficult to implement in FDD systems and static environments.
PLA • Be realized based on physical-layer attributes, mainly including channel fingerprints and RF fingerprints. • Channel fingerprints are derived from the unique characteristics of signals affected by the propagation environment, such as channel state information (CSI), path loss, and delay spread. Channel fingerprints are unique in both time and space, and have short-term stability. • RF fingerprints originates from the manufacturing differences of equipment hardware, such as In-phased/Quadrature (I/Q) imbalance, carrier frequency offset (CFO), and phase noise. These differences lead to a unique “hardware fingerprint” embedded in the transmitted signal, which has hardware uniqueness and long-term stability. Even the devices of the same model have slight differences in their RF fingerprints, which are not easy to be covered up by software modification or environmental interference. •  Pros: Endogenous security. The endogenous security characteristics of channel fingerprints and RF fingerprints stem from the uniqueness and non-replicability of their physical properties, which are directly determined by the physical-layer characteristics of the communication system itself, without relying on external encryption algorithms or key management. • Pros: Lightweight. Firstly, it circumvents upper-layer signaling processes. Additionally, because access points have obtained CSI of legitimate users in channel estimation stages, the calculation overhead is further reduced. • Pros: High compatibility. Even if devices can not decode each other’s upper-layer signaling, as long as they share the same physical-layer communication standard, they can still decode the physical-layer bit stream and extract the fingerprints needed for PLA, thus realizing cross-device and cross-system authentication, which provides strong support for interoperability and compatibility between devices.
@@ II-B1 What is GAI?
GAI represents a category of AI technology aimed at synthesizing new data instances that adhere to a specified probability distribution. Unlike conventional DAI optimized for classification or regression tasks, GAI excels at learning potential structural patterns within training data, enabling autonomous generation of diverse content such as images, audio, and text. Its strength resides in modeling the distribution of high-dimensional data, facilitating the generation of realistic synthetic datasets, and augmenting sparse training samples. These capabilities make GAI particularly valuable in areas where data is scarce or labeling costs are too high, where DAI struggles.
@@ II-B2 How to GAI?
TABLE II: The outline of the issues faced in the field of PLA, the limitations of Discriminative Artificial Intelligence (DAI) technologies in tackling these difficulties, and the promising capabilities of Generative AI (GAI) to boost performance within these contexts, where insufficient fingerprint data occur due to the unpredictable interference of environmental complexity and the inherent physical constraints of hardware measurement systems; environment noises and inferences appear due to path loss variations caused by obstacle obstruction, and rapid signal intensity fluctuations resulting from multipath effect; fingerprint data is perturbed by adversarial attacks and backdoor attacks; and complex tasks include multi-user authentication tasks, identifying new users joining the system, and realizing multiuser authentication and attack detection.
@@
Insufficient fingerprint data • It heavily relies on a large number of labeled fingerprint samples to define classification boundaries, and depends on upper-layer authentication protocols or manual operations to assign identity labels of fingerprint samples. • This strict supervised learning mode encounters constraints in annotation costs. When fingerprint samples are scarce, it is prone to overfitting or inadequate feature representation. • The potential value of unlabeled fingerprint samples remains underutilized, resulting in a waste of data resources. • It learns distributions of fingerprints rather than directly constructing mapping relationships between labels and fingerprints. • It not only significantly expands the size of training fingerprint sets but also allows for the controlled generation of fingerprint variants simulating different conditions, thereby enhancing the PLA model’s adaptability to data scarcity. • It takes the fingerprint data distribution itself as the learning object, rather than relying on identity labels, thus fully utilizing the potential value of unlabeled data [11].
Environment noises and inferences • It tends to treat noise as a static interference factor independent of fingerprint features, which restricts the model to constructing classification boundaries only under preset SNR conditions. • It prioritizes classification accuracy as its core optimization objective. The learning process often focuses on capturing the most easily distinguishable superficial features in noisy environments, but struggles to deeply analyze the complex nonlinear interactions between noise and fingerprints. • It constructs a joint probability space encompassing both fingerprints and noises. Thus, it explicitly expresses interaction through a mathematical structure, revealing the deep-seated patterns of noise’s intrinsic influence on fingerprint features [12]. • By incorporating SNR as a conditional input, It not only breaks the dependence of generated fingerprint samples on specific SNR conditions but also fundamentally improves the generalization when faced with varying SNRs [1].
Perturbations in fingerprint data • It is often limited to optimizing the classification boundaries of known clean fingerprint data, lacking proactive strategies to defend against adversarial attacks. • It typically relies on static rules or preset thresholds for authentication decisions, a rigid decision-making mechanism that is difficult to cope with evolving interference patterns. • It is highly sensitive to minor changes in input fingerprints. • GAN possesses an inherent defensive advantage due to its adversarial training nature [13]. • It cleverly formulates PLA as probability inference problems by learning the underlying probability distribution of fingerprints. • When input signals are disturbed, it can reconstruct the features of the original signal based on the learned fingerprint distribution, thereby eliminating the impact of perturbation components.
Complex tasks • When faced with highly similar or significantly overlapping fingerprint features among multiple users, the classification boundaries are forced to be compressed into narrow regions, leading to ambiguity in decision-making. • It requires re-adjusting classification boundaries to adapt to the updated user categories. • It typically requires separately constructing multiuser authentication models and attack detection models, and achieving collaboration between them through series or parallel connections. • It builds an exclusive probability distribution model for the fingerprint features of each user individually, rather than forcing fingerprint features of all users into the same space. • It only needs to train the generator branch for updated users without modifying existing user modules or the global model structure. This “plug-and-play” feature greatly reduces the maintenance complexity of the PLA system. • It integrates multiuser authentication and attack detection into a single-stage task through joint probability modeling [5].
@@ II-C What PLA Issues Can Benefit from GAI?
As depicted in Table II, the limitations of DAI and GAI’s potential in PLA are presented as follows.
  • Insufficient Fingerprint Data: Given the unpredictable interference caused by environmental complexity and the inherent physical constraints of hardware measurement systems, it is often difficult to collect sufficient high-precision fingerprint samples in actual communication scenarios [3]. Traditional DAI methods have limitations in relying on a large number of fingerprint samples, relying on upper-layer authentication protocols or manual labeling of fingerprints, and low utilization of unlabeled fingerprint data. In contrast, GAI has potential in augmenting fingerprint data, learning fingerprint data distributions through latent variables, learning unlabeled fingerprint data, and continuously optimizing its model parameters and structure.
@@
  • Perturbations in Fingerprint Data: Faced with emerging intelligent threats such as adversarial attacks [13] and backdoor attacks [2], traditional DAI methods have limitations, lacking defense against well-designed adversarial perturbations, decision making based on fixed rules, and vulnerability and sensitivity to small disturbances. In contrast, GAI has the potential to defend against adversarial attacks through adversarial training, eliminate interference through fingerprint reconstruction, and make flexible inferences, judgments, and decisions based on learned fingerprint probability distributions.
  • Complex Tasks: For complex tasks such as multiuser authentication and identifying updated user categories, traditional DAI methods have limitations in fuzzy classification boundaries caused by overlapping fingerprint feature spaces of users and require retraining the PLA model to adapt to the updated user categories. In contrast, GAI has the potential to generate independent fingerprint distributions for each user to avoid feature overlap and expand to updated users without retraining the PLA model.
Figure 1: Illustration of GAI-enhanced methods for PLA, where GAI models include Generative Adversarial Network (GAN), Variational Autoencoder (VAE), Generative Diffusion Model (GDM), and Transformer; and PLA issues are classified according to fingerprint collection, model training, and performance optimization phases. The fingerprint collection stage provides fingerprint dataset for model training stage, and the performance of the trained PLA model is further enhanced in the performance optimization stage.
TABLE III: The use of GAI in PLA.
@@
Fingerprint Transmission - Encoders and decoders for joint source-channel coding Channel adaptive de-noising transmission -
@@ III GAI-enhanced Methods for PLA
As illustrated in Figure 1, the implementation of GAI-based PLA necessitates a self-evolving system comprising three stages: fingerprint collection, model training, and performance optimization.
  • During the fingerprint collection phase, GAI empowers fingerprint estimation, fingerprint augmentation, fingerprint filling, and fingerprint transmission to improve accuracy, enrich diversity, ensure completeness, and enhance transmission efficiency, respectively, thereby providing reliable fingerprint samples for training PLA models.
  • In the model training phase, GAI-driven training strategies are tailored to specific task requirements, such as unsupervised attack detection, multi-user authentication, and efficient training and deployment, to develop PLA models.
  • During the performance optimization phase, the trained PLA model undergoes further enhancement through GAI to achieve advanced capabilities, including improving flexibility for eliminated legitimate users, enhancing robust defense against intelligent attacks, and boosting adaptability across diverse scenarios.
Table III further compares how representative GAI models address PLA issues from each of these phases, with detailed descriptions provided below.
@@ III-A1 Fingerprint Estimation
Channel estimation is a critical technique for acquiring channel fingerprints. DAI-based estimation methods require long pilot sequences to achieve satisfactory performance and struggle in high path loss scenarios. In contrast, GAI can generate synthetic training fingerprints with distributions similar to real training samples, thereby providing effective support for fingerprint estimation models. For instance, the variational inference process in VAEs approximates the maximum likelihood process of fingerprint estimation. Similarly, GDMs can learn score functions of posterior fingerprint distributions and employ posterior sampling during reverse processes to recover fingerprint samples. Additionally, Zhang et al. [10] applied the discriminator in GANs to learn and extract intrinsic features of time-varying channels in high-speed railway systems, while the generator determines implicit mapping functions of training fingerprint data to enhance estimation performance.

### v3

@@ Abstract
Recently, Artificial Intelligence (AI)-driven Physical-Layer Authentication (PLA), which focuses on achieving endogenous security and intelligent identity authentication, has attracted considerable interest. When compared with Discriminative AI (DAI), Generative AI (GAI) offers several advantages, such as fingerprint augmentation, reconstruction, and denoising. Inspired by these innovations, this paper provides a systematic exploration of GAI’s integration with PLA. We commence with a concise review of identity authentication techniques and GAI models. Then, we contrast the limitations of DAI with the potential of GAI in addressing PLA challenges. Specifically, we introduce a structured taxonomy for GAI-enhanced PLA methodologies, encompassing three key stages: fingerprint collection, model training, and performance optimization within the PLA pipeline. Furthermore, we propose a novel PLA framework based on GAI and channel extrapolation for dynamic environments. To demonstrate GAI’s efficacy in enhancing PLA robustness, we implement a case study using the Generative Diffusion Model (GDM). Finally, we outline potential future research directions for GAI-based PLA.
@@ I Introduction
Traditional Discriminative AI (DAI) utilizes deep neural networks to learn features from large-scale fingerprint datasets. However, it relies on extensively labeled fingerprints for training. Fortunately, the emergence of Generative AI (GAI) has mitigated these limitations. Instead of learning class boundaries, GAI learns the probability distribution of training fingerprints to generate new fingerprint samples [14], which is expected to be applied to Internet of Things (IoT) [11, 7], Industrial IoT (IIoT) [9], Internet of Vehicles (IoV) [12], and satellite communication systems [5]. For instance, Meng et al. [9] proposed a Hierarchical Variational Autoencoder (HVAE)-based PLA scheme to detect spoofing attackers at the edge of IIoT. Besides, Liu et al. [7] proposed a Bidirectional Encoder Representation from Transformers (BERT)-based lightweight PLA framework to enhance the zero-trust edge IoT security. Additionally, Jiang et al. [5] combined convolutional neural networks, Generative Adversarial Networks (GANs), and Bayesian meta-learning, to realize security authentication of the satellite component in automatic identification systems. Furthermore, Chi et al. [1] combined Generative Diffusion Models (GDMs) and Transformers to learn the time, frequency and complex-valued domains of RF fingerprints. Nevertheless, research on GAI-based PLA is still in its early phases. The full potential of applying GAI to PLA has yet to be comprehensively investigated. Motivated by this, we present a systematic investigation into GAI’s integration with PLA. The primary contributions are outlined as:
  • We analyze the issues of traditional DAI solved by GAI in PLA, including insufficient fingerprint data, environmental noises and interferences, perturbations in fingerprint data, and complex tasks.
  • We propose a structured taxonomy for GAI-enhanced PLA methods, delineating three stages within the PLA pipeline: fingerprint collection, model training, and performance optimization. The fingerprint collection stage encompasses fingerprint estimation, fingerprint augmentation, fingerprint filling, and fingerprint transmission. The model training stage includes attack detection, multiuser authentication, and efficient training and deployment. The performance optimization stage focuses on the flexibility for eliminated legitimate users,defense against intelligent attacks, and adaptation for different scenarios.
  • We propose a channel extrapolation-based PLA framework integrating GAI, addressing the challenges faced by conventional PLA schemes. We further present a novel case study leveraging GDMs, to show the superiority of GAI in enhancing the robustness of PLA.
  • We outline a future research roadmap for GAI-based PLA across five critical dimensions: high-reliable fingerprints, efficient PLA, generalization enhancement, model security, and industrialization.
@@ II-A1 Why Identity Authentication?
  • Preventing Unauthorized Access: Identity authentication verifies the authenticity of communicating parties, blocks unauthorized devices at the source, and establishes a trusted communication link [9].
  • Safeguarding Data Security: Identity authentication acts as a trust anchor for encryption systems, ensuring that keys are exclusively held by legitimate parties, thereby preserving the confidentiality and integrity of data transmission [7].
@@ II-A2 How to Identity Authentication?
Four representative identity authentication techniques are compared as follows. Cryptography-based upper-layer authentication is mainly used in mobile communication systems, but relies on the limited computation power assumptions of attackers; Blockchain-based authentication is based on a multi-layer framework and a fault-tolerant mechanism, but has high delay and low throughput; Physical-layer-key-based authentication relies on the wireless channel reciprocity, but is difficult to implement in frequency division duplexing systems and static environments. In contrast, PLA leverages physical-layer attributes, including channel fingerprints and RF fingerprints, to realize endogenous security, lightweight, and high compatibility.
@@ II-B1 What is GAI?
Unlike conventional DAI optimized for classification or regression tasks, GAI excels at learning potential structural patterns within training data, enabling autonomous generation of diverse content such as images, audio, and text. Its strength resides in modeling the distribution of high-dimensional data, facilitating the generation of realistic synthetic datasets, and augmenting sparse training samples.
@@ II-B2 How to GAI?
TABLE I: The outline of the issues faced in the field of PLA, the limitations of DAI in tackling these difficulties, and the promising capabilities of GAI to boost performance within these contexts.
@@
Insufficient fingerprint data • It relies on a large number of labeled fingerprint samples. • This strict supervised learning mode increases annotation costs. • The potential value of unlabeled fingerprint samples remains underutilized. • It directly learns distributions of fingerprints. • It allows for the controlled generation of fingerprint variants simulating different conditions. • It fully utilizes the potential value of unlabeled data [15].
Environment noises and inferences • It constructs classification boundaries only under preset SNR conditions. • It struggles to deeply analyze the complex nonlinear interactions between noise and fingerprints. • It reveals the deep-seated patterns of noise’s intrinsic influence on fingerprint features by joint probability modeling [1]. • It breaks the dependence of generated fingerprint samples on specific SNR conditions [11].
Perturbations in fingerprint data • It lacks proactive strategies to defend against adversarial attacks. • It relies on static rules for authentication decisions. • It is highly sensitive to minor changes in input fingerprints. • GAN possesses an inherent defensive advantage due to its adversarial training nature [10]. • It cleverly formulates PLA as probability inference problems. • It can eliminate the impact of perturbation components through the learned fingerprint distribution.
Complex tasks • When faced with highly similar fingerprint features among multiple users, it is difficult to distinguish them. • It requires retraining to adapt to the updated user categories. • It typically requires separately constructing models for different tasks. • It builds an exclusive probability distribution model for the fingerprint features of each user. • It introduces fine-tuning methods to adapt to the updated users without retraining the model. • It integrates multi-tasks through joint probability modeling [3].
@@ II-C What PLA Issues Can Benefit from GAI?
As depicted in Table I, the limitations of DAI and GAI’s potential in PLA are presented as follows.
  • Insufficient Fingerprint Data: Given the unpredictable interference caused by environmental complexity and the inherent physical constraints of hardware measurement systems, it is often difficult to collect sufficient high-precision fingerprint samples in actual communication scenarios [9]. Traditional DAI methods have limitations in reliance on a large number of fingerprint samples, reliance on upper-layer authentication protocols or manual labeling of fingerprints, and low utilization of unlabeled fingerprint data. In contrast, GAI has potential in augmenting fingerprint data, learning fingerprint data distributions through latent variables, learning unlabeled fingerprint data, and continuously optimizing its model parameters and structure.
@@
  • Perturbations in Fingerprint Data: Faced with emerging intelligent threats such as adversarial attacks [10], traditional DAI methods have limitations, lacking defense against well-designed adversarial perturbations, decision making based on fixed rules, and vulnerability and sensitivity to small disturbances. In contrast, GAI has the potential to defend against adversarial attacks through adversarial training, eliminate interference through fingerprint reconstruction, and make flexible inferences, judgments, and decisions based on learned fingerprint probability distributions.
  • Complex Tasks: For complex tasks such as multiuser authentication and identifying the updated user categories, traditional DAI methods have limitations in fuzzy classification boundaries caused by overlapping fingerprint feature spaces of users and require retraining the PLA model. In contrast, GAI has the potential to generate independent fingerprint distributions for each user to avoid feature overlap and expand to the updated users without retraining the PLA model.
Figure 1: Illustration of GAI-enhanced methods for PLA, where Part (A) denotes representative GAI models, including Generative Adversarial Network (GAN), Variational Autoencoder (VAE), Generative Diffusion Model (GDM), and Transformer, Part (B) denotes GAI for fingerprint collection phase, including fingerprint estimation, fingerprint augmentation, fingerprint filling, and fingerprint transmission, Part (C) denotes GAI for PLA model training, including attack detection, multiuser authentication, and channel knowledge map, and Part (D) denotes GAI for PLA performance optimization, including flexibility for eliminated legitimate users, defense against intelligent attacks, and adaptation for different scenarios.
TABLE II: The applications of GAI models in PLA.
@@
Fingerprint Transmission Reconstructing transmitted signal using generators Encoders and decoders for joint source-channel coding Channel adaptive de-noising transmission Attention-based semantic encoding and decoding
@@ III GAI-enhanced Methods for PLA
As illustrated in Figure 1, we sort out which stages of PLA can be enhanced by GAI, including fingerprint collection, authentication model training, and authentication performance optimization.
@@ III-A1 Fingerprint Estimation
DAI-based estimation methods require long pilot sequences to achieve satisfactory performance and struggle in high path loss scenarios. In contrast, GAI can generate synthetic training fingerprints with distributions similar to real training samples, thereby providing effective support for fingerprint estimation. Zhang et al. [12] applied the discriminator in GANs to learn and extract intrinsic features of time-varying channels in high-speed railway systems, while the generator determines implicit mapping functions of training fingerprint data to enhance estimation performance.